I WANT ARITHEMATIC PROGRESIONS IMPORTANT FOR SCHOOL SHUFFLING TEST

Grade:10

1 Answers

SHAIK HAFEEZUL KAREEM
8 years ago
Types of Progression: Progressions are of 3 types. They are
1. Arithmetic Progression (A.P.)
2. Geometric Progression (G.P)
3. Harmonic Progression (H.P)
1. Arithmetic Progression: A progression in which, except the first term, each term is
obtained by adding a fixed number to the term immediately preceding it is called arithmetic
progression.
♦ The fixed number is called "common difference" and it is denoted by "d".
♦ The terms of the progression are denoted by t 1 , t 2 , t 3 , t 4 , ......., first term t 1 is also denoted
by "a".
♦ Common difference (d) = t 2 – t 1 = t 3 – t 2 = t 4 – t 3 ........
d = t 2 − t 1
♦ In 21, 16, 11, 6,....... common difference is –5.
d = t 2 – t 1
= 16 – 21
= –5
♦ n th term or general term of A.P. t n = a + ( n − 1 ) d
plese approve my answer i collect all the information about AP